OpenStudy (anonymous):

QUESTION 6 Given A(6, -17), B(-6, 17), C(4, 9), D(-4, 4), the midpoint of which segment is at the origin? AB BC AC AD 1 points QUESTION 7 Which is the equation of the line passing through the points A(3, -3) and B(-4, 11)? y = -6x + 3 y = -2x + 3 y = 9x -3 y = 4x - 9

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

second part: I apply this formula: \[\frac{ y-y _{1} }{ y _{2}-y _{1} }=\frac{ x-x _{1} }{ x _{2}-x _{1} }\] where (x_1,y_1) and (x_2, y_2) are your points. so:

OpenStudy (michele_laino):

\[\frac{ y-(-3) }{ 11-(-3) }=\frac{ x-3 }{ -4-3 }\]

OpenStudy (michele_laino):

then, after some simplification, we have: \[y+3=-2(x-3)=-2x+6\] finally: \[y=-2x+3\]

OpenStudy (anonymous):

ok and #1?

OpenStudy (michele_laino):

first part: as usually if we apply the same formula, to find midpoint of a segment given its endpoints, we have: midpoint AB=(0,0), namely it is the origin of our coordinate system, midpoint BC=(-1,13), midpoint AC=(5,-4), midpoint AD=(1,-13/2)

OpenStudy (michele_laino):